OpenStudy (anonymous):

At constant volume, if the Kelvin temperature of a gas is doubled, the pressure of the gas is halved. A. True B. False

OpenStudy (anonymous):

@aprehan :)

OpenStudy (anonymous):

I'm going to let you think through this after I give you some helpful information. Since R = PV/nT, if the Temperature is increased (specifically doubled), what do you think must happen to the Pressure?

OpenStudy (anonymous):

if my downstairs increases, what must happen to the upstairs in order to keep R constant?

OpenStudy (anonymous):

The same thing ? So its false?

OpenStudy (anonymous):

if I have 10 = 5/2, then lets say i increase my downstairs 2 into 5, then that means my upstairs of 5 must be increased to 50 in order for my ratio to still be 10 (b/c 50/5 is 10)

OpenStudy (anonymous):

Yes it's false

OpenStudy (anonymous):

The pressure would also increase (specifically double)

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (sweetburger):

say this is your original 1L/1K=k you now double it to 2L what must the new Temperature in Kelvins be to equal the same constant